40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1171     Entry time: Tue Dec 2 16:12:18 2014
Author: Dmass 
Type: Update 
Category: Roadmap 
Subject: The week's measurements 

The following measurements are being done this week:

 

Photothermal transfer function at 120K

Wilcoxen seismometer (accelerometer?) added to setup

Cold seismic noise (boiling noise)

Cold temp noise

Cold beat w/ Weiner subtraction of PDH** / intensity / seismic

**PDH error signal f-domain Weiner subtraction @ low gain regime

ELOG V3.1.3-